OpenStudy (anonymous):

For order of operations what comes first Parentheses or Exponents? Woud 2(-1)^2 be 4 or 2 or -4?!

OpenStudy (anonymous):

my calculator says 2. I say it lies.

OpenStudy (anonymous):

P E M D A S

myininaya (myininaya):

2(-1)^2=2(1)=2

OpenStudy (anonymous):

thanks! Calculator wasn't lying I guess! I tried doing 2*-1 =-2^2 and it was lying then!

myininaya (myininaya):

2*(-1) does not equal -2^2